# Survey and Experiments on Mental Disorder Detection via Social Media: From Large Language Models and RAG to Agents

### v3

@@ Abstract
Mental disorders represent a critical global health challenge, and social media is increasingly viewed as a vital resource for real-time digital phenotyping and intervention. Large Language Models (LLMs) offer stronger semantic understanding and reasoning than traditional deep learning, but their use in high‑stakes clinical settings is limited by hallucinations and the lack of persistent memory. However, existing literature has not sufficiently investigated how advanced enhancement techniques, specifically Retrieval-Augmented Generation (RAG) and Agentic systems, can address these reliability and reasoning limitations. Here, we systematically survey the evolving landscape of LLM-based methods for social media mental disorder analysis, spanning standard pre-trained language models, RAG to mitigate hallucinations and contextual gaps, and agentic systems for autonomous reasoning and multi-step intervention. We organize existing work by technical paradigm and clinical target, extending beyond common internalizing disorders to include psychotic disorders and externalizing behaviors. Additionally, the paper comprehensively evaluates the performance of LLMs, including the impact of RAG, across various tasks. This work establishes a unified benchmark for the field, paving the way for the development of trustworthy, autonomous AI systems that can deliver precise and explainable mental health support.
@@ I Introduction
As a result, social media has emerged as an important source[naslund2020social, chen2021social], providing real-time spontaneous expressions that are less susceptible to the reporting biases found in traditional questionnaires. According to [social-app-report], the number of active users on social media platforms is immense, and these platforms have become the most important venues for social interaction and information dissemination, which in turn generates a massive, real-time volume of user-generated content (such as posting updates, sharing personal status, and complaining about various issues) that reflects not only users’ daily lives but also their emotional fluctuations and psychological states[zhang2022natural]. Furthermore, users can interact with others through retweets, comments, likes, and other actions, enabling them to express their personal feelings and thoughts more comprehensively, generating a vast amount of real-time social information of significant value.
However, social media data is inherently unstructured, multimodal, and characterized by noise, sparsity, and high velocity [abkenar2021big, gao2022modeling], which collectively present significant challenges. Therefore, exploring how to develop advanced data mining and knowledge discovery techniques to effectively utilize this complex social media data for mental health monitoring and prediction has become a key focus for future research.
While traditional machine learning, deep learning, and early pre-trained models (e.g., BERT) laid the foundation for text analysis, they often fall short in reasoning capabilities and interpretability required for mental health tasks. With the rapid development of natural language processing, large language models (LLMs) [hadi2023survey, li2024survey] have bridged this gap. These models transcend simple classification by offering a deep understanding of syntactic and semantic contexts [mudrik2024exploring, omar2024applications]. They enable not only the detection of mental disorder symptoms from user posts but also the generation of diagnostic explanations based on rich pre-trained knowledge. Despite these advancements, applying LLMs in high-stakes domains remains challenging. Without specialized medical grounding, LLMs can suffer from ‘hallucinations’, generating unfaithful or fabricated responses [zhang2023large, huang2025survey], which is unacceptable for mental disorder detection. Moreover, LLMs lack persistent memory and explicit planning, making them insufficient for executing complex, multi-step clinical decision workflows.
@@
While existing surveys [wongkoblap2017researching, skaik2020using, harrigian2020state, lejeune2022use, malhotra2022deep, garg2023mental, di2023methodologies, owen2024ai, shah2024mental, omar2024exploring, hua2024large, hua2024applying, ke2024exploring, guo2024large, lawrence2024opportunities] have explored various approaches and challenges in detecting multiple mental health problems, a significant gap remains: none has investigated the role of RAG techniques and agentic systems. This oversight leaves unexplored how RAG integration enhances factual reliability through external knowledge retrieval, and how agentic systems introduce a paradigm shift towards autonomous reasoning and persistent memory, enabling the orchestration of complex diagnostic workflows in the noisy social media environment. Additionally, many current surveys [hua2024large, hua2024applying, ke2024exploring, guo2024large, lawrence2024opportunities] focus solely on the application of LLMs in the mental health field without specializing in social media, where individuals frequently express mental disorders [naslund2020social]. Neglecting social media data, which provides real-time insights for early intervention and long-term data to model disease progression, limits our understanding of how LLMs can be effectively applied to these dynamic platforms.
@@ V-C2 Overall Performance
Model Metric CSSRS-Suicide SDCNL STRD DepSeverity Dreaddit SWMH-Anxiety SWMH-Bipolar GPT-4o $\Delta$ W-F1 $\uparrow$ +0.024 $\uparrow$ +0.005 $\downarrow$ -0.027 $\downarrow$ -0.068 $\uparrow$ +0.062 $\uparrow$ +0.030 $\uparrow$ +0.028 $\Delta$ B-Accuracy $\uparrow$ +0.028 $\uparrow$ +0.004 $\downarrow$ -0.028 $\downarrow$ -0.015 $\uparrow$ +0.058 $\uparrow$ +0.008 $\uparrow$ +0.010 Qwen3-32B $\Delta$ W-F1 $\downarrow$ -0.032 $\uparrow$ +0.024 $\downarrow$ -0.026 $\downarrow$ -0.012 $\downarrow$ -0.048 $\downarrow$ -0.010 $\uparrow$ +0.006 $\Delta$ B-Accuracy $\downarrow$ -0.002 $\uparrow$ +0.021 $\downarrow$ -0.003 $\downarrow$ -0.006 $\downarrow$ -0.037 $\uparrow$ +0.056 $\uparrow$ +0.007 DeepSeek-R1 $\Delta$ W-F1 $\downarrow$ -0.047 $\downarrow$ -0.010 $\downarrow$ -0.012 $\downarrow$ -0.056 $\uparrow$ +0.018 $\downarrow$ -0.009 $\uparrow$ +0.032 $\Delta$ B-Accuracy $\downarrow$ -0.015 $\downarrow$ -0.09 $\uparrow$ +0.003 $\downarrow$ -0.013 $\uparrow$ +0.018 $\uparrow$ +0.012 $\uparrow$ +0.021 DeepSeek-V3.1 $\Delta$ W-F1 $\uparrow$ +0.039 $\uparrow$ +0.004 $\downarrow$ -0.026 $\downarrow$ -0.065 $\uparrow$ +0.040 $\downarrow$ -0.005 $\uparrow$ +0.004 $\Delta$ B-Accuracy $\uparrow$ +0.036 $\uparrow$ +0.002 $\downarrow$ -0.023 $\downarrow$ -0.022 $\uparrow$ +0.034 $\uparrow$ +0.014 $\uparrow$ +0.025 Llama-3.3-70B $\Delta$ W-F1 $\downarrow$ -0.018 $\uparrow$ +0.076 $\downarrow$ -0.054 $\downarrow$ -0.028 $\uparrow$ +0.063 $\downarrow$ -0.015 $\uparrow$ +0.003 $\Delta$ B-Accuracy $\uparrow$ +0.012 $\uparrow$ +0.052 $\downarrow$ -0.038 $\downarrow$ -0.023 $\uparrow$ +0.042 $\downarrow$ -0.014 $\uparrow$ +0.005 Llama-3.1-405B $\Delta$ W-F1 $\uparrow$ +0.025 $\downarrow$ -0.022 $\uparrow$ +0.042 $\downarrow$ -0.05 $\uparrow$ +0.041 $\downarrow$ -0.026 $\downarrow$ -0.042 $\Delta$ B-Accuracy $\uparrow$ +0.014 $\uparrow$ +0.003 $\uparrow$ +0.04 $\downarrow$ -0.007 $\uparrow$ +0.035 $\downarrow$ -0.003 $\downarrow$ -0.014 Llama-3.1-70B $\Delta$ W-F1 $\downarrow$ -0.011 $\uparrow$ +0.058 $\downarrow$ -0.079 $\downarrow$ -0.117 $\uparrow$ +0.059 $\downarrow$ -0.028 $\uparrow$ +0.021 $\Delta$ B-Accuracy $\uparrow$ +0.053 $\uparrow$ +0.031 $\downarrow$ -0.115 $\downarrow$ -0.045 $\uparrow$ +0.047 $\downarrow$ -0.075 $\uparrow$ +0.024 Note. Each cell reports the change in weighted F1 ($\Delta$ W-F1) or balanced accuracy ($\Delta$ B-Accuracy) when moving from the zero-shot to the few-shot setting (few-shot minus zero-shot). Positive values ($\uparrow$ +) indicate improvements from few-shot prompting, whereas negative values ($\downarrow$ -) indicate performance drops. This table is calculated between the zero-shot and the few-shot sections of Table III.

### v4

@@ Abstract
Mental disorders represent a critical global health challenge, and social media is increasingly viewed as a vital resource for real-time digital phenotyping and intervention. To leverage this data, large language models (LLMs) have been introduced, offering stronger semantic understanding and reasoning than traditional deep learning, thereby enhancing the explainability of detection results. Despite the growing prominence of LLMs in this field, there is a scarcity of scholarly works that systematically synthesize how advanced enhancement techniques, specifically Retrieval-Augmented Generation (RAG) and Agentic systems, can be utilized to address these reliability and reasoning limitations. Here, we systematically survey the evolving landscape of LLM-based methods for social media mental disorder analysis, spanning standard pre-trained language models, RAG to mitigate hallucinations and contextual gaps, and agentic systems for autonomous reasoning and multi-step intervention. We organize existing work by technical paradigm and clinical target, extending beyond common internalizing disorders to include psychotic disorders and externalizing behaviors. Additionally, the paper comprehensively evaluates the performance of LLMs, including the impact of RAG, across various tasks. This work establishes a unified benchmark for the field, paving the way for the development of trustworthy, autonomous AI systems that can deliver precise and explainable mental health support.
@@ I Introduction
As a result, social media has emerged as an important source [naslund2020social, chen2021social], providing large-scale, real-time and relatively spontaneous expressions that reflects not only users’ daily lives but also their emotional fluctuations and psychological states, which are less susceptible to the reporting biases found in traditional questionnaires. With an immense active user base [social-app-report], these platforms generate massive data that capture both daily behaviors and emotional and psychological signals. According to Zhang et al. [zhang2022natural], among the 399 studies included in their review, social media text dominates as the primary data source (81%), far exceeding interviews (7%) and electronic health records (6%).
However, social media data is inherently unstructured and characterized by semantic complexity and label sparsity [abkenar2021big, gao2022modeling]. Effectively extracting reliable mental health signals from such high-entropy environments necessitates systems capable of deep semantic interpretation and complex reasoning. Without these capabilities, detection models remain dependent on extensive labeled datasets and capture only surface-level statistical correlations, often missing the nuanced psychological intent hidden within the noise.
While traditional machine learning, deep learning, and early pre-trained language models (e.g., BERT) laid the foundation for text analysis, they often fall short in meeting these reasoning and interpretability demands. With the rapid development of natural language processing, large language models (LLMs) [hadi2023survey, li2024survey] have bridged this gap. With their pre-trained world knowledge and superior reasoning capabilities, LLMs can effectively interpret these subtle mental health signals even in low-resource settings. These models transcend simple classification by offering a deep understanding of syntactic and semantic contexts [mudrik2024exploring, omar2024applications]. They enable not only the detection of mental disorder symptoms from user posts but also the generation of diagnostic explanations based on rich pre-trained knowledge. Despite these advancements, applying LLMs in high-stakes domains remains challenging. Without specialized medical grounding, LLMs can suffer from ‘hallucinations’, generating unfaithful or fabricated responses [zhang2023large, huang2025survey], which is unacceptable for mental disorder detection. Moreover, LLMs lack persistent memory and explicit planning, making them insufficient for executing complex, multi-step clinical decision workflows.
@@
While existing surveys [wongkoblap2017researching, skaik2020using, harrigian2020state, lejeune2022use, malhotra2022deep, garg2023mental, di2023methodologies, owen2024ai, shah2024mental, omar2024exploring, hua2024large, hua2024applying, ke2024exploring, guo2024large, lawrence2024opportunities] have explored different approaches and challenges in detecting mental disorders, a significant gap remains: none have investigated the role of RAG techniques and agentic systems. This oversight leaves unexplored how RAG integration enhances factual reliability through external knowledge retrieval, and how agentic systems introduce a paradigm shift towards autonomous reasoning and persistent memory, enabling the orchestration of complex diagnostic workflows in the noisy social media environment. Additionally, many current surveys [hua2024large, hua2024applying, ke2024exploring, guo2024large, lawrence2024opportunities] primarily focus on the general application of LLMs in the mental health field, with less emphasis placed on social media contexts, where individuals frequently express mental disorder symptoms [naslund2020social]. Neglecting social media data, which provides real-time insights for early intervention and long-term data to model disease progression, limits our understanding of how LLMs can be effectively applied to these dynamic platforms.
@@ V-C2 Overall Performance
Model Metric CSSRS-Suicide SDCNL STRD DepSeverity Dreaddit SWMH-Anxiety SWMH-Bipolar GPT-4o $\Delta$ W-F1 $\uparrow$ +0.024 $\uparrow$ +0.005 $\downarrow$ -0.027 $\downarrow$ -0.068 $\uparrow$ +0.062 $\downarrow$ -0.030 $\uparrow$ +0.028 $\Delta$ B-Accuracy $\uparrow$ +0.028 $\uparrow$ +0.004 $\downarrow$ -0.028 $\downarrow$ -0.015 $\uparrow$ +0.058 $\uparrow$ +0.008 $\uparrow$ +0.010 Qwen3-32B $\Delta$ W-F1 $\downarrow$ -0.032 $\uparrow$ +0.024 $\downarrow$ -0.026 $\downarrow$ -0.012 $\downarrow$ -0.048 $\downarrow$ -0.010 $\uparrow$ +0.006 $\Delta$ B-Accuracy $\downarrow$ -0.002 $\uparrow$ +0.021 $\downarrow$ -0.003 $\downarrow$ -0.006 $\downarrow$ -0.037 $\uparrow$ +0.056 $\uparrow$ +0.007 DeepSeek-R1 $\Delta$ W-F1 $\downarrow$ -0.047 $\uparrow$ +0.010 $\downarrow$ -0.012 $\downarrow$ -0.056 $\uparrow$ +0.018 $\downarrow$ -0.009 $\uparrow$ +0.032 $\Delta$ B-Accuracy $\downarrow$ -0.015 $\uparrow$ +0.009 $\uparrow$ +0.003 $\downarrow$ -0.013 $\uparrow$ +0.018 $\uparrow$ +0.012 $\uparrow$ +0.021 DeepSeek-V3.1 $\Delta$ W-F1 $\uparrow$ +0.039 $\uparrow$ +0.004 $\downarrow$ -0.026 $\downarrow$ -0.065 $\uparrow$ +0.040 $\downarrow$ -0.005 $\uparrow$ +0.040 $\Delta$ B-Accuracy $\uparrow$ +0.036 $\uparrow$ +0.002 $\downarrow$ -0.023 $\downarrow$ -0.022 $\uparrow$ +0.034 $\uparrow$ +0.014 $\uparrow$ +0.025 Llama-3.3-70B $\Delta$ W-F1 $\downarrow$ -0.018 $\uparrow$ +0.076 $\downarrow$ -0.054 $\downarrow$ -0.028 $\uparrow$ +0.053 $\downarrow$ -0.015 $\uparrow$ +0.003 $\Delta$ B-Accuracy $\uparrow$ +0.012 $\uparrow$ +0.052 $\downarrow$ -0.038 $\downarrow$ -0.023 $\uparrow$ +0.042 $\downarrow$ -0.014 $\uparrow$ +0.005 Llama-3.1-405B $\Delta$ W-F1 $\uparrow$ +0.025 $\downarrow$ -0.022 $\uparrow$ +0.042 $\downarrow$ -0.050 $\uparrow$ +0.041 $\downarrow$ -0.026 $\downarrow$ -0.042 $\Delta$ B-Accuracy $\uparrow$ +0.014 $\uparrow$ +0.003 $\uparrow$ +0.040 $\downarrow$ -0.007 $\uparrow$ +0.035 $\downarrow$ -0.003 $\downarrow$ -0.014 Llama-3.1-70B $\Delta$ W-F1 $\downarrow$ -0.011 $\uparrow$ +0.058 $\downarrow$ -0.115 $\downarrow$ -0.117 $\uparrow$ +0.059 $\downarrow$ -0.028 $\uparrow$ +0.021 $\Delta$ B-Accuracy $\uparrow$ +0.053 $\uparrow$ +0.031 $\downarrow$ -0.079 $\downarrow$ -0.045 $\uparrow$ +0.047 $\downarrow$ -0.075 $\uparrow$ +0.024 Note. Each cell reports the change in weighted F1 ($\Delta$ W-F1) or balanced accuracy ($\Delta$ B-Accuracy) when moving from the zero-shot to the few-shot setting (few-shot minus zero-shot). Positive values ($\uparrow$ +) indicate improvements from few-shot prompting, whereas negative values ($\downarrow$ -) indicate performance drops. This table is calculated between the zero-shot and the few-shot sections of Table III.